### v1

# Gas-induced perturbations on the gravitational wave in-spiral of live post-Newtonian LISA massive black hole binaries

Abstract
We investigate the effect of dynamically coupling gas torques with gravitational wave (GW) emission during the orbital evolution of an equal-mass massive black hole binary (MBHB). We perform hydrodynamical simulations of eccentric MBHBs with total mass $M=10^{6}~{}{\rm M}_{\sun}$ embedded in a prograde locally isothermal circumbinary disk (CBD). We evolve the binary from $53$ to $30$ Schwarzschild radii separations using up to 2.5 post-Newtonian (PN) corrections to the binary dynamics, which allow us to follow the GW-driven in-spiral. For the first time, we report the measurement of gas torques onto a live binary a few years before the merger, with and without concurrent GW radiation. We also identify and measure a novel GW-gas coupling term in the in-spiral rate that makes gas effects an order of magnitude stronger than the gas-only contribution. We show that the evolution rate ($\dot{a}$) of the MBHB can be neatly expressed as the sum of the GW rate ($\dot{a}_{\rm GW}$), the pure gas-driven rate ($\dot{a}_{\rm gas}$), and their cross-term $\propto\dot{a}_{\rm GW}\dot{a}_{\rm gas}$. The source-frame gas-induced dephasing in the GW waveform is equivalent to losing $\sim 0.5$ GW cycles over the expected $\sim 1700$ cycles in a vacuum, which LISA should detect at redshift $z=1$. We also propose a phenomenological model that captures the essence of simulations and can be used to perform Bayesian inference. Our results show how GWs alone can be used to probe the astrophysical properties of CBDs and have important implications on multi-messenger strategies aimed at studying the environments of MBHBs.
accretion, accretion disks — black hole physics — gravitational waves — hydrodynamics — relativistic processes — (galaxies:) quasars: supermassive black holes
1 Introduction
The recent adoption of LISA (Amaro-Seoane et al., 2017; Colpi et al., 2024) and the development of TianQin (Li et al., 2024) and Taiji (Gong et al., 2021) will provide a powerful opportunity to detect gravitational waves (GWs) from coalescing near-equal mass massive black hole binaries (MBHBs) with masses $\sim 10^{4}$-$10^{7}~{}{\rm M}_{\sun}$. LISA can potentially detect MBHBs up to redshifts $z\lesssim 20$ and with high (e.g. $\lesssim 10^{3}$) signal-to-noise ratios (SNRs; Amaro-Seoane et al., 2017). MBHBs are a by-product of galaxy mergers (Begelman et al., 1980). When two galaxies merge, the massive black holes (MBHs) hosted in their centre are expected to reach the centre of the remnant galaxy owing to the dynamical friction mechanism and form a bound binary at a pc scale. This binary can further proceed towards merger through the interaction with surrounding stars and gas until GWs are strong enough to take over and drive the binary to coalescence (Amaro-Seoane et al., 2023).
While interactions with stars in a tri-axial potential (e.g. Quinlan, 1996; Preto et al., 2011; Khan et al., 2011) as well as a third close-by MBH (Blaes et al., 2002; Hoffman & Loeb, 2007; Bonetti et al., 2019) can lead to a MBHB merger, they are relatively slow and rare mechanisms, respectively. In particular, sinking timescales of MBHBs due to three body encounters with stars can exceed a Gyr in the low density environments of stellar-disk dominated galaxies (Khan et al., 2018), the typical hosts of MBHs below $10^{6}~{}{\rm M}_{\sun}$, that fall in the mass range accessible by LISA. On the other hand, when the host galaxies are gas-rich, and have circumnuclear gas disks, then MBHBs can sink efficiently below pc separations (Mayer, 2013; Souza Lima et al., 2020). At even smaller separations, below $\sim 0.1$ pc, the MBHB would carve a cavity, which results in the formation of a co-planar geometrically thin (Shakura & Sunyaev, 1973) circumbinary disk (CBD; Escala et al. 2004; Cuadra et al. 2009; D’Orazio et al. 2016), allowing the binary to coalesce in less than $100$ Myr (Haiman et al., 2009). Over the past few years, the interaction between a binary and its circumbinary disk has been studied extensively for various system parameters and thermodynamics assumptions using several different numerical hydrodynamical (HD) simulations (see e.g. Duffell et al., 2024). The general consensus is that circular nearly equal-mass binaries do undergo out-spiral in relatively thick CBDs while their in-spiral is aided by relatively thin (i.e. aspect ratio $H/R\lesssim 0.03$) CBDs (Tiede et al., 2020; Franchini et al., 2021, 2022).
Since MBHBs observable by LISA are likely to reside in gaseous environments (see, e.g. Mangiagli et al., 2022), it is important to study the effect of gas on the orbital evolution of the MBHB when it enters the LISA band. The first attempt in this direction was performed by several groups (Garg et al., 2022; Tiede et al., 2024; Zwick et al., 2024; Garg et al., 2024a) who measured the gas-induced dephasing in the LISA band by simply linearly adding the gas-driven evolution rate, computed in post-processing from 2D HD fixed binary orbit simulations, to the GW in-spiral rate. However, the scales considered in those numerical works are close to sub-pc, where GWs are still too weak to drive significant binary evolution. Furthermore, by adding the two contribution linearly, gas-induced dephasing studies might have ignored possible coupling between gas torques and GW-driven evolution, due to the lack of HD simulations where the two effects are naturally coupled together and the binary evolves under both processes at the same time.
Recently, Franchini et al. (2024) simulated an eccentric, live (Franchini et al., 2023), equal-mass $10^{6}~{}{\rm M}_{\sun}$ MBHB embedded in a prograde $100{\rm M}_{\sun}$ CBD by dynamically modeling the binary in-spiral with post-Newtonian (PN) corrections up to $2.5$ order. They evolved the system for the final years of in-spiral, including the merger and post merger phase, to quantify possible electromagnetic (EM) counterparts. In this work, we use the same setup to simulate the same binary but now embedded in a lighter $5~{}{\rm M}_{\sun}$ disk to properly investigate how gas perturbs the binary evolution rate. We then quantify for the first time the effect of gas-induced perturbations on waveforms using a live binary whose dynamics is computed using PN corrections, thus including the interplay between energy and angular momentum change caused by both GW radiation and gas torques. With this simulation setup, the in-spiral is concurrently determined by GWs and gas, as opposed to co-adding the two effects in post-processing as previously done in the literature, allowing us to robustly quantify the gas-induced dephasing in the GW waveform and its detectability by LISA.
2 Numerical setup
Following the approach used in Franchini et al. (2024), we model the binary using two equal-mass sink particles (Bate et al., 1995) that represent two Schwarzschild MBHs with total mass $M=10^{6}~{}{\rm M}_{\sun}$ , setting the respective sink radius to the innermost stable circular orbit (ISCO). We set the MBHB initial semi-major axis (SMA) to $a=53.2$ Schwarzschild radii ($r_{s}$) and eccentricity to $e=0.29$. Note that we start from the initial condition of the thin (i.e. aspect ratio $H/R=0.03$) disk simulation in Franchini et al. (2024), which originated from a circular equal-mass binary evolved for $1000$ binary orbits by Franchini et al. (2022). During the first $1000$ orbits the eccentricity of the simulated live binary increases to $e=0.29$ as a result of the interaction with the CBD. We here assume the disk to have a much smaller mass $M_{\rm d}=5{\rm M}_{\sun}$.
We follow the evolution of the binary driven by both gas torques and PN corrections up to 2.5PN order using the code gizmo (Hopkins, 2015) until the binary reaches $=30\,r_{\rm s}$ in separation in our GW+gas simulation, i.e. for $\sim 1700$ GW cycles. The implementation of the PN corrections to the binary dynamics follows the equations in Blanchet (2014). We include both conservative $1$PN and $2$PN terms, and radiative $2.5$PN terms. The latter term generates the GW emission and leads to the decrease in binary SMA ($\dot{a}_{\rm GW}$) and eccentricity only due to GWs. In order to integrate the 2.5PN equations, we implemented an intermediate predictor step to update the particle velocities at the end of the time step, accounting for the PN corrections, and re-enforcing the numerical stability of the integration algorithm. Our approach is similar to the one outlined in Sect. 6.2 of Liptai & Price (2019), except that we use a predictor-corrector approach instead of implementing the implicit kick-drift-kick one. In the initial setup the disk extended from $2a$ to $10a$. However during the first 1000 binary orbits, the cavity becomes eccentric and the inner edge increases to $\sim 3.5a$.
We also run two other simulations with the same setup. First, without the dissipative $2.5$PN term to consider the binary evolution only due to the presence of the CBD. We run it for $1000$ orbits to see any appreciable changes in the orbital quantities. This allows us to infer the SMA evolution rate only due to the gaseous disk ($\dot{a}_{\rm gas}$) and therefore to disentangle its effect from the putative GW-gas cross term considered later. We then also run a simulation without the CBD in order to self-consistently obtain the evolution due to the PN terms only. This allows us to mitigate the numerical error in the integration when we compute the difference between the simulation with concurrent evolution under GWs and gas, and the one with PN terms only.
2.1 Post-processing analysis
The CBD typically affects the binary evolution by exerting both a gravitational torque ($T_{\rm grav}$) and an accretion torque ($T_{\rm acc}$). The first, just due to gravity, is essentially driven by any asymmetry in the gaseous flow while the latter is instead induced by the accretion of gas particles onto either MBHs. In particular, the accretion of gas alters not only the mass but also the angular momentum of the binary (see Franchini et al. 2021 for detailed calculations).
We can then express the overall gas effect in terms of a single simulation-calibrated parameter (Garg et al., 2022; Duffell et al., 2024)
$\xi=\frac{T_{\rm grav}+T_{\rm acc}}{\dot{M}a^{2}\Omega},$ (1)
where $\dot{M}\equiv 0.02~{}{\rm f}_{\rm Edd}(M/10^{6}{\rm M}_{\sun}){({\rm M}_{\sun}/\rm yr)}$ is the accretion rate onto the binary for our assumed $\epsilon=0.1$ radiative efficiency, $\Omega\equiv\sqrt{GM/a^{3}}$ is the binary orbital angular frequency, and $\dot{M}a^{2}\Omega$ is the viscous torque (Lin & Papaloizou, 1986). $\xi$ depends sensitively on the binary and disk parameters, in particular on the binary mass ratio $q$, disk shape and temperature, and may also depends upon the assumed equation of state. Previous sub-pc fixed binary orbit 2D Newtonian simulation predicted $|\xi|\sim 3$ (Dittmann & Ryan, 2022) for $H/R\sim 0.03$.
We can therefore write down the SMA evolution rate due to gas torques for a quasi-circular binary as (Garg et al., 2024a)
$\dot{a}_{\rm gas}=8\xi\frac{\dot{M}_{z}}{M_{z}}a=\xi{\rm f}_{\rm Edd}\frac{4}{25{\rm Myr}}a,$ (2)
where we infer both $\xi$ and ${\rm f}_{\rm Edd}$ from gas-only simulation to properly isolate any GW-gas coupling in the total SMA rate in the GW+gas simulation. Moreover, since the evolution only due to gas torques changes the SMA by a small quantity even after $1000$ orbits, we extrapolate the results to smaller separations, i.e. until $30r_{s}$.
When the MBHB is dynamically evolving under both GWs and gas, there can be a coupled evolution that manifests in the form of a cross term $\dot{a}_{\rm GWgas}$, whose exact functional form is unknown. Therefore, the total SMA evolution rate can be written down as
$\dot{a}_{\rm GW+gas}=\dot{a}_{\rm GW}+\dot{a}_{\rm gas}+\dot{a}_{\rm GWgas},$ (3)
where both $\dot{a}_{\rm gas}$ and $\dot{a}_{\rm GWgas}$ can be thought of as perturbative corrections to $\dot{a}_{\rm GW}$ in the final years of in-spiral.
Assuming the stationary phase approximation (Cutler & Flanagan, 1994), where gas affects the phase but not the amplitude of the GW, we compute the dephasing in the time-domain GW waveform, caused by the presence of the CBD as
$\delta\phi_{\rm gas}=\Phi(\dot{a}_{\rm GW+gas})-\Phi(\dot{a}_{\rm GW}),$ (4)
where negative (positive) dephasing imply that gas speed-up (slow-down) the binary in-spiral. Here $\Phi(\dot{a})$ is the total accumulated GW phase for a given evolution rate $\dot{a}$ between an initial ($a_{i}$) and a final ($a_{f}$) SMAs, and it can be expressed in radians as
$\Phi(\dot{a})=2\pi\int_{a_{\rm i}}^{a_{\rm f}}{\rm d}a\frac{f}{\dot{a}},$ (5)
where $f\equiv\Omega/\pi$ is the GW frequency, that is twice the orbital frequency ($\Omega/2\pi$), under the assumption of quasi-circular binaries.
In the next section, we show the results together with the interpretation of our simulations.
3 Results
We show disk morphologies and associated column densities at three SMAs – $a=50r_{s}$, $a=40r_{s}$, and $a=30r_{s}$ – in Fig. 1 for the GW+gas simulation.111disk morphology for gas-only simulation is similar to the left panel of Fig. 1. The gas morphology is similar, as expected, to the simulations presented in Franchini et al. (2024). We can clearly see the over-density at the cavity edge, i.e. the lump (Shi et al., 2012), precessing around the binary.
We measure the accretion rate in both our simulations and find a mean Eddington ratio of ${\rm f}_{\rm Edd}\approx 0.92$ for GW+gas simulation and ${\rm f}_{\rm Edd}\approx 1.20$ for the gas-only case. The slightly higher accretion rate in the gas-only case is a natural consequence of the fact that the binary does not start decoupling from the disk because of the lack of GW radiation, therefore the gas can flow inside the cavity and keep feeding the MBHB more effectively.
Figure 1: Column density ($\Sigma$) plots at three SMAs: $50r_{s}$ (left panel), $40r_{s}$ (middle panel), and $30r_{s}$ (right panel) for the binary evolution under both GW and gas. Here $\Sigma$ varies between $\sim 10^{3.5}$-$10^{6}$ g/cm2. While both the binary (green dots) and the cavity shrinks with time, cavity is becoming more eccentric also.
We then compute both gravitational and accretion torques directly from the simulations. We find that the magnitude of the accretion torque is $T_{\rm acc}\sim 10^{-2}T_{\rm grav}$. We show the evolution of the parameter $\xi$ in Fig. 2 for GW+gas simulations as a function of SMA together with the average values of $\xi$ for both GW+gas ($\bar{\xi}_{\rm GW+gas}\approx-24$) and gas-only ($\bar{\xi}_{\rm gas}\approx-36$) cases. Unsurprisingly, the mean gas effect is stronger when the binary is only evolving due to the gas torque. The parameter $\xi$ is indeed decreasing in magnitude as the binary is in-spiraling further, since the gas effects are becoming weaker. Since the value of $\xi$ does not change significantly with time in the gas-only simulation, the approximation we make in the parametrization of Eq. (1) is reasonable. It is worth noticing that for the computed gas-induced dephasing $\delta\phi_{\rm gas}$ (see Section 3.1), most of the contribution is expected to originate from the early in-spiral phase, where the mean value of $\xi$ is higher.
We also note that the high-frequency or sub-orbital fluctuations in Fig. 2 around the mean value may be measurable on their own, as suggested by analytical studies (Zwick et al., 2021, 2024). However, we focus on the secular trend in this work.
Figure 2: Gas torques onto the binary in terms of $\xi$ as a function of the SMA for GW+gas simulation (light blue lines). We show average $\xi$ values for GW+gas (solid blue line; $\bar{\xi}_{\rm GW+gas}\approx-24$) and gas-only (dashed red line; $\bar{\xi}_{\rm gas}\approx-36$) cases, respectively.
Once we subtract both $\dot{a}_{\rm GW}$ and $\dot{a}_{\rm gas}$ from the total SMA $\dot{a}_{\rm GW+gas}$ rate inferred from the full GW+gas simulation, we find a non-negligible residual, which is the GW-gas coupling term introduced in Eq. (3). We propose to express this cross-term simply as
$\dot{a}_{\rm GWgas}\sim\zeta\dot{a}_{\rm GW}\dot{a}_{\rm gas},$ (6)
where $\zeta$ is also a simulation-calibrated parameter and we show its evolution in Fig. 3 as function of SMA. We find the average value of the cross-term to be $\approx-7.9\times 10^{-5}$. The value of $\zeta$ is also decreasing as the binary is shrinking, similar to the evolution of $\xi$ in Fig. 2. The almost constant evolution of $\zeta$, marginally weakened by the in-spiral, may imply that our assumption to express $\dot{a}_{\rm GWgas}\propto\dot{a}_{\rm GW}\dot{a}_{\rm gas}$ is not inaccurate.
Figure 3: GW-gas coupling coefficient as a function of SMA (light blue lines) with its mean value $\bar{\zeta}=-7.9\times 10^{-5}$ (solid blue line).
3.1 Gas-induced dephasing and LISA observability
In this section, we compute the gas-induced dephasing $\delta\phi_{\rm gas}$ in Eq. 4 in the source-frame between $=53.2~{}r_{s}$ and $=30~{}r_{s}$ with the initial eccentricity $=0.29$ by using $\dot{a}_{\rm GW}$ inferred from the simulations and then considering the total SMA evolution rate $\dot{a}_{\rm GW+gas}$ in three different ways.
  • First of all directly from the GW+gas simulations, which leads to a dephasing $\delta\phi_{\rm gas}=-3.21$ radians.
  • The second estimate of the dephasing is performed by neglecting the GW-gas cross term, i.e. assuming $\dot{a}_{\rm GW+gas}=\dot{a}_{\rm GW}+\dot{a}_{\rm gas}$, where $\dot{a}_{\rm gas}$ is computed from Eq. 2 with the gas-only simulation-calibrated parameters $\xi=-36$ and ${\rm f}_{\rm Edd}=1.2$. This approach gives $\delta\phi_{\rm gas}=-0.21$ radians.
  • Lastly, we include the cross-term using the parameterization in Eq. 6 to have $\dot{a}_{\rm GW+gas}=\dot{a}_{\rm GW}+\dot{a}_{\rm gas}+\zeta\dot{a}_{\rm GW}\dot{a}_{\rm gas}$ with $\zeta=-7.9\times 10^{-5}$. This results in $\delta\phi_{\rm gas}=-3.63$ radians.
The dephasing we obtain neglecting the GW-gas cross term (i.e. assuming $\zeta=0$) is an order of magnitude smaller than the value inferred with the non-zero cross term, so we caution that the GW-gas coupling might have a crucial role in the evolution of MBHBs in gaseous environments. Furthermore, the fact that the $\delta\phi_{\rm gas}$ computed directly from the GW+gas simulation is almost the same as the dephasing computed by parametrizing the gas effect through the (simulation-calibrated) $\xi$ and $\zeta$ parameters, provides a further consistency check of our model. The few percent difference between the two values might originate from the approximation made in assuming $\xi$ and $\zeta$ to be constants with time.
We consider the evolution of our $10^{6}~{}{\rm M}_{\sun}$ total mass eccentric MBHB at redshift $z=1$ for one year of LISA observation. We therefore take an initial SMA of $a_{i}=40r^{z=1}_{s}$ and an initial eccentricity $e=0.2$, as the binary has partially circularized from its initial separation. We find the dephasing for ${\rm f}_{\rm Edd}=1.2$, $\xi=-36$, and $\zeta=-7.9\times 10^{-5}$ between $a_{i}=40r^{z=1}_{s}$ and $a_{f}=30r^{z=1}_{s}$ to be $\delta\phi_{\rm gas}=-1.11$ radians, equivalent to losing $\sim 0.2$ GW cycles over the expected $\sim 850$ cycles from a vacuum eccentric binary evolving just under GWs. Since LISA should observe this event with SNR1300similar-toabsent1300$\sim$ 1300∼ 1300 (Garg et al., 2024d) and we typically need a given dephasing to be higher than 8$\sim$ 8/∼ 8 /SNR (Derdzinski et al., 2021; Garg et al., 2022) to be detectable, we should be able to observe the computed $\delta\phi_{\rm gas}$.
3.1.1 Analytical dephasing estimates
Since we would ultimately like to obtain a parametrized model for the GW+gas inspiral phase, we now use the orbital-averaged GW-driven inspiral rate at $2.5$PN order from Peters (1964), i.e. $\dot{a}^{\rm Peters}_{\rm GW}$, for the GW only driven inspiral rate. Moreover, we also analytically construct the total SMA rate using the simulation-calibrated parameters $\xi$, ${\rm f}_{\rm Edd}$, and $\zeta$ to have
$\dot{a}^{\rm analytical}_{\rm GW+gas}=\dot{a}^{\rm Peters}_{\rm GW}+\dot{a}_{\rm gas}+\zeta\dot{a}^{\rm Peters}_{\rm GW}\dot{a}_{\rm gas}.$ (7)
If we repeat the dephasing computation done in § 3.1, then we get $\delta\phi_{\rm gas}=-3.75$ radians. This is a few percent higher than the value calculated using only the simulations data. One possible reason is the inclusion of 1PN and 2PN terms in the simulations which, despite being conservative terms, lead to the precession of the MBHB eccentric orbit effectively altering its in-spiral rate. Another source of uncertainty in the dephasing estimate might reside in numerical errors. We have however compared our GW-only simulation with a numerical integration of the full (i.e. up to 2.5 order) PN evolution of the binary using an 8th order Runge-Kutta integrator finding negligible differences within the separations range we explored.
Even though we have only performed simulations for one set of binary-disk parameters, we can compute the expected dephasing for different values of the $\xi$ and $\zeta$ parameters, which reflect the use of different system parameters in the simulations, through the analytical prescription in Eq. 7. In Fig. 4, we show $\delta\phi_{\rm gas}$ for equal-mass circular MBHBs accreting at ${\rm f}_{\rm Edd}=1$ for a one-year LISA observation until respective ISCO with varying $\xi\in[-1,-100]$, $\zeta\in[0,2\times 10^{-4}]$, $M\in\{10^{5},10^{6}\}~{}{\rm M}_{\sun}$, and $z\in\{1,2\}$ to bracket reasonable parameters space. If we do not consider the GW-gas cross-term then the dephasing cannot be detectable by LISA for $M=10^{6}~{}{\rm M}_{\sun}$ MBHB at both redshifts for $|\xi|\lesssim 1$. The stronger dephasing with non-zero $\zeta$ again signifies the importance of considering the GW-gas cross-term in having observable gas effects on the GW waveform. Moreover, similar $\delta\phi_{\rm gas}$ for different combinations of mass and redshift is possibly due to the cross-term dependence on $\dot{a}^{\rm Peters}_{\rm GW}$.
Figure 4: The gas-induced dephasing in the GW waveform, $\delta\phi_{\rm gas}$ in radians, over a one-year LISA observation until ISCO as a function of the simulation-calibrated fudge factors $\xi$ (represents gas torques) and $\zeta$ (represents GW-gas cross-term) for equal-mass MBHBs accreting at ${\rm f}_{\rm Edd}=1$. In all four panels, we vary $\xi$ from $-1$ to $-100$ and $\zeta$ from 00 to $2\times 10^{-4}$. In the top two panels, we consider systems at redshift $z=2$ and in the bottom two panels we assume $z=1$. In the left two panels, we study a MBHB with source-frame total mass $M=10^{5}~{}{\rm M}_{\sun}$ and in the right panels we have $M=10^{6}~{}{\rm M}_{\sun}$. We also draw cyan lines to indicate the regions of non-observable dephasing by LISA, i.e. $\delta\phi_{\rm gas}<8/$SNR.
A perturbative correction to $\dot{a}_{\rm GW}$ with a functional form $\propto a^{\rm ng}\dot{a}_{\rm GW}$ manifests at $-n_{g}$ relative order in the GW phase (Garg et al., 2024a). We have $n_{g}=4$ and $n_{g}=1$ for $\dot{a}_{\rm gas}$ and $\dot{a}_{\rm GWgas}$, respectively. Hence, an instantaneous frequency-depending dephasing enters the overall binary phase at relative $-4$PN order due to evolution just under gas and at $-1$PN relative order due to a GW-gas coupling.
Lastly, if we do not introduce a GW-gas coupling and try to explain the residual $(\dot{a}_{\rm GW+gas}-\dot{a}_{\rm GW})$ by using the model for $\dot{a}_{\rm gas}$ in Eq. 2, then the mean value of the parameter $\xi$ is $\sim-360$. This implies that even a dephasing at $-4$PN relative order should be observable for our studied MBHB with or without concurrent eccentricity measurement (Garg et al., 2024a) in a realistic Bayesian data analysis setup, where all other binary parameters are also varying and a complete LISA response is considered. However, the GW-gas coupling is even more stronger than this, since it enters the GW phase at $-1$PN relative order, therefore causing a larger dephasing.
4 Discussion and Conclusion
We studied the interaction of an equal mass MBHB with its surrounding CBD during the late in-spiral stage, with and without concurrent GW emission, using 3D hydrodynamical simulations with a live binary. This approach provided us with the first direct measurement of how surrounding gas torques the binary when its in-spiral is already governed by GW emission, by means of the estimate of the $\xi$ parameter in Eq. (1). We find that $\xi$ is $\mathcal{O}(10)$ stronger than the previously explored scale-free/sub-pc regime. However, we caution that the comparison may not be fair as previous simulations carried out for the larger separation regime are predominantly 2D, assume a fixed binary orbit modelled under Newtonian dynamics, and compute the effect of the energy and angular momentum loss by GW radiation in post-processing (Tang et al., 2018). In particular, Duffell et al. (2024) showed that 3D calculations give different magnitude torques compared to 2D and Franchini et al. (2023) argued that fixing the binary orbit leads again to a different gravitational torque. Moreover, we find that, consistently with our expectations, $\xi$ is slightly weaker for GW+gas simulations ($\xi_{\rm GW+gas}\sim-24$) with respect to gas-only study ($\xi_{\rm GW+gas}\sim-36$), as shown in Fig. 2. This is expected since the binary is decoupling from the gas in the GW+gas simulation and therefore the effect of gas weakens with time.
Most importantly, we show that our results can be understood by introducing a novel coupling term between the energy loss by GWs and that due to gas in the orbital in-spiral rate, which we model through the product between the rate of change of the SMA due to GWs only and that due to gas only. The amplitude of this coupling term is pinpointed by a newly introduced simulation-calibrated parameter $\zeta$, whose evolution is shown in Fig. 3 with an average value $\sim-7.9\times 10^{-5}$. This implies richer gas features in the GW waveform than previously expected. While gas torques still induce a correction to the GW phase at the usual $-4$PN relative order (Garg et al., 2024a), our modeling of GW-gas cross-term enters the phase at $-1$PN relative order, leading to stronger phase shift by $\mathcal{O}(10)$ and higher likelihood of detection of gas perturbations with LISA. The $-1$PN relative order term may also allow us to break the degeneracy with concurrent eccentricity measurement using LISA (Garg et al., 2024a). Moreover, having a GW-gas coupling may change the expectation how the binary parameters in the LISA band are correlated with their values at sub-pc scales (Garg et al., 2024c). Lastly, ignoring gas effects while modeling MBHBs may lead to stronger apparent violations of general relativity than previously anticipated (Garg et al., 2024b) due to the new $-1$PN relative order term.
In this work, we studied one set of binary-disk parameters while in the future we plan on exploring the parameter space to understand how both $\xi$ and $\zeta$ change. This exercise may also allow us to analytically model both simulation-calibrated parameters $\xi$ and $\zeta$. For instance, if the binary is embedded in a relatively thick disk ($H/R=0.1$), we can expect to also have a significant contribution from mini-disks (i.e. disks around each binary component) that will introduce their own likely positive torque measurement (D’Orazio & Duffell, 2021; Franchini et al., 2022; Siwek et al., 2023; Tiede & D’Orazio, 2023) and probably a different sign for $\zeta$. Moreover, different numerical studies may also allow to check the robustness of our phenomenological model.
Lastly, we also find that, in the gas-only simulation, the eccentricity is linearly decreasing with the SMA. This needs to be further investigated in both sub-pc and near-merger regimes, and we plan to do so in the future.
It is to be noted that we employ PN corrections as an approximation since we do not run general relativistic magnetohydrodynamical (GRMHD) simulations. Current GRMHD simulations only study the binary evolution just a few days before merger (Gutiérrez et al., 2022; Avara et al., 2023) to integrate only a few orbits because of the prohibitive computational cost. Therefore, since the majority of the gas-driven effects on the binary in-spiral occurs at separations $a\gtrsim 30~{}r_{s}$, our approach is currently the best available method to investigate the orbital dephasing due to the presence of the gas and where PN corrections are may be adequate.
A possible caveat in our work is that our binary is moderately eccentric ($\sim 0.3$) just a few years before the merger, which will require the eccentricity to be extremely high when GWs take over at milli-pc scales. However, since $e\sim 0.3$ arises naturally from our initial condition requiring a steady-state disk before setting the physical scale of $a\sim 53r_{s}$, the only truly realistic way to initialize the system is to evolve the binary starting from a much wider separation. This, however, would increase the computational cost dramatically. We plan to investigate alternative procedures in the setup of the simulations in order to reduce the computational cost and mitigate this issue in the future.
In summary, our results can facilitate the modeling of gas effects perturbing GW waveforms, which in turn will allow to better quantify how effectively LISA can place constraints on the environment of MBHBs, eventually opening the pathway for more informed synergies between GWs and EM observations. Furthermore, our work, being the first of its kind with PN dynamics and a live binary in 3D, while still assuming a simple isothermal equation of state, provides a starting point for future hydrodynamical studies with additional physics, including, for example, more realistic thermodynamics.
Acknowledgements
MG acknowledge support from the Swiss National Science Foundation (SNSF) under the grant 200020_192092. The authors also acknowledge use of the NumPy (Harris et al., 2020) and Matplotlib (Hunter, 2007).

### v3

@@
 We perform hydrodynamical simulations of eccentric MBHBs with total mass $M=10^{6}~{\rm M}_{\sun}$ embedded in a prograde locally isothermal circumbinary disk (CBD).
-We evolve the binary from $53$ to $30$ Schwarzschild radii separations using up to 2.5 post-Newtonian (PN) corrections to the binary dynamics, which allow us to follow the GW-driven in-spiral.
+We evolve the binary from $55$ to $49$ Schwarzschild radii separations using up to 2.5 post-Newtonian (PN) corrections to the binary dynamics, which allow us to follow the GW-driven in-spiral.
 For the first time, we report the measurement of gas torques onto a live binary a few years before the merger, with and without concurrent GW radiation.
-We also identify and measure a novel GW-gas coupling term in the in-spiral rate that makes gas effects an order of magnitude stronger than the gas-only contribution.
-We show that the evolution rate ($\dot{a}$) of the MBHB can be neatly expressed as the sum of the GW rate ($\dot{a}_{\rm GW}$), the pure gas-driven rate ($\dot{a}_{\rm gas}$), and their cross-term $\propto\dot{a}_{\rm GW}\dot{a}_{\rm gas}$.
-The source-frame gas-induced dephasing in the GW waveform is equivalent to losing $\sim 0.5$ GW cycles over the expected $\sim 1700$ cycles in a vacuum, which LISA should detect at redshift $z=1$.
-We also propose a phenomenological model that captures the essence of simulations and can be used to perform Bayesian inference.
-Our results show how GWs alone can be used to probe the astrophysical properties of CBDs and have important implications on multi-messenger strategies aimed at studying the environments of MBHBs.
+We also report the gas-induced orbital dephasing $\delta\phi_{\rm orb}\sim-0.007$ rad over $278$ orbital cycles that is likely driven mainly by disk-induced precession and LISA should be able to detect it at redshift $z=1$.
+Our results show how GWs alone can be used to probe the astrophysical properties of CBDs and have important implications for multi-messenger strategies aimed at studying the environments of MBHBs.
 accretion, accretion disks — black hole physics — gravitational waves — hydrodynamics — relativistic processes — (galaxies:) quasars: supermassive black holes
 1 Introduction
-The recent adoption of LISA (Amaro-Seoane et al., 2017; Colpi et al., 2024) and the development of TianQin (Li et al., 2024) and Taiji (Gong et al., 2021) will provide a powerful opportunity to detect gravitational waves (GWs) from coalescing near-equal mass massive black hole binaries (MBHBs) with masses $\sim 10^{4}$-$10^{7}~{}{\rm M}_{\sun}$.
+The recent adoption of LISA (Amaro-Seoane et al., 2017; Colpi et al., 2024) and the development of TianQin (Li et al., 2025) and Taiji (Gong et al., 2021) will provide a powerful opportunity to detect gravitational waves (GWs) from coalescing near-equal mass massive black hole binaries (MBHBs) with masses $\sim 10^{4}$-$10^{7}~{\rm M}_{\sun}$.
 LISA can potentially detect MBHBs up to redshifts $z\lesssim 20$ and with high (e.g.
@@
 MBHBs are a by-product of galaxy mergers (Begelman et al., 1980).
-When two galaxies merge, the massive black holes (MBHs) hosted in their centre are expected to reach the centre of the remnant galaxy owing to the dynamical friction mechanism and form a bound binary at a pc scale.
+When two galaxies merge, the massive black holes (MBHs) hosted in their centre are expected to reach the centre of the remnant galaxy owing to the dynamical friction mechanism and form a bound binary at pc scales.
 This binary can further proceed towards merger through the interaction with surrounding stars and gas until GWs are strong enough to take over and drive the binary to coalescence (Amaro-Seoane et al., 2023).
@@
 On the other hand, when the host galaxies are gas-rich, and have circumnuclear gas disks, then MBHBs can sink efficiently below pc separations (Mayer, 2013; Souza Lima et al., 2020).
-At even smaller separations, below $\sim 0.1$ pc, the MBHB would carve a cavity, which results in the formation of a co-planar geometrically thin (Shakura & Sunyaev, 1973) circumbinary disk (CBD; Escala et al.
+At separations below $\sim 0.1$ pc, the gas dragged by the two MBHs is expected to settle in a common circumbinary disk (CBD; Escala et al.
 2004; Cuadra et al.
 2009; D’Orazio et al.
-2016), allowing the binary to coalesce in less than $100$ Myr (Haiman et al., 2009).
-Over the past few years, the interaction between a binary and its circumbinary disk has been studied extensively for various system parameters and thermodynamics assumptions using several different numerical hydrodynamical (HD) simulations (see e.g.
+2016), whose torques remove angular momentum from the binary allowing it to coalesce in less than $100$ Myr (Haiman et al., 2009).
+The binary potential will open a cavity inside the CBD whose size depends on the disk properties (Artymowicz & Lubow, 1994, 1996).
+Over the past few years, the interaction between a binary and its circumbinary disk has been studied extensively for various system parameters and thermodynamics assumptions using different numerical hydrodynamical (HD) simulations (see e.g.
 Duffell et al., 2024).
@@
 Mangiagli et al., 2022), it is important to study the effect of gas on the orbital evolution of the MBHB when it enters the LISA band.
-The first attempt in this direction was performed by several groups (Garg et al., 2022; Tiede et al., 2024; Zwick et al., 2024; Garg et al., 2024a) who measured the gas-induced dephasing in the LISA band by simply linearly adding the gas-driven evolution rate, computed in post-processing from 2D HD fixed binary orbit simulations, to the GW in-spiral rate.
+The first attempt in this direction was performed by several groups (Garg et al., 2022; Tiede et al., 2024; Zwick et al., 2024; Garg et al., 2024a, c, b; Dittmann et al., 2023) who measured the gas-induced dephasing in the LISA band by simply linearly adding the gas-driven evolution rate, computed in post-processing from 2D HD fixed binary orbit simulations, to the GW in-spiral rate.
 However, the scales considered in those numerical works are close to sub-pc, where GWs are still too weak to drive significant binary evolution.
@@
 Recently, Franchini et al.
-(2024) simulated an eccentric, live (Franchini et al., 2023), equal-mass $10^{6}~{}{\rm M}_{\sun}$ MBHB embedded in a prograde $100{\rm M}_{\sun}$ CBD by dynamically modeling the binary in-spiral with post-Newtonian (PN) corrections up to $2.5$ order.
-They evolved the system for the final years of in-spiral, including the merger and post merger phase, to quantify possible electromagnetic (EM) counterparts.
+(2024) simulated an eccentric, live (Franchini et al., 2023), equal-mass $10^{6}~{\rm M}_{\sun}$ MBHB embedded in a prograde $100~{\rm M}_{\sun}$ CBD by dynamically modeling the binary in-spiral with post-Newtonian (PN) corrections up to $2.5$ order.
+They evolved the system for the final years of in-spiral, including the merger and post-merger phase, to quantify possible electromagnetic (EM) counterparts.
 In this work, we use the same setup to simulate the same binary but now embedded in a lighter $5~{\rm M}_{\sun}$ disk to properly investigate how gas perturbs the binary evolution rate.
@@
 2 Numerical setup
-Following the approach used in Franchini et al.
-(2024), we model the binary using two equal-mass sink particles (Bate et al., 1995) that represent two Schwarzschild MBHs with total mass $M=10^{6}~{}{\rm M}_{\sun}$ , setting the respective sink radius to the innermost stable circular orbit (ISCO).
-We set the MBHB initial semi-major axis (SMA) to $a=53.2$ Schwarzschild radii ($r_{s}$) and eccentricity to $e=0.29$.
-Note that we start from the initial condition of the thin (i.e.
-aspect ratio $H/R=0.03$) disk simulation in Franchini et al.
-(2024), which originated from a circular equal-mass binary evolved for $1000$ binary orbits by Franchini et al.
-(2022).
-During the first $1000$ orbits the eccentricity of the simulated live binary increases to $e=0.29$ as a result of the interaction with the CBD.
-We here assume the disk to have a much smaller mass $M_{\rm d}=5{\rm M}_{\sun}$.
-We follow the evolution of the binary driven by both gas torques and PN corrections up to 2.5PN order using the code gizmo (Hopkins, 2015) until the binary reaches $=30\,r_{\rm s}$ in separation in our GW+gas simulation, i.e.
-for $\sim 1700$ GW cycles.
+[…]
+We explore three resolutions for our gas+PN2.5 setup by increasing the number of splitting levels in the hyper-Lagrangian refinement.
+We quantify the resolution in terms of inter-particle spacing $\Delta x$ at $R=3a$.
+We label these simulations low-resolution (LR) with $\Delta x[3a]={1.31~r_{s}}$, mid-resolution (MR) with $\Delta x[3a]={1.07~r_{s}}$, and high-resolution (HR) with $\Delta x[3a]={0.66~r_{s}}$.
+We run at least $100$ orbits for each simulations in order to perform a meaningful resolution study, which we report in Appendix A.
+We find that the gravitational torque exerted by the disk onto the binary in the MR simulation is already converged and therefore we consider the MR run as our fiducial setup.
+Unless stated otherwise, the following results have been inferred using our MR run.
+We follow the evolution of the binary driven by both gas torques and PN corrections up to 2.5PN order using the code gizmo (Hopkins, 2015) until the binary reaches $48.9\,r_{\rm s}$ in separation in our gas+PN2.5 simulation, i.e.
+for $278$ initial binary orbits or $\sim 556$ GW cycles.
 The implementation of the PN corrections to the binary dynamics follows the equations in Blanchet (2014).
-We include both conservative $1$PN and $2$PN terms, and radiative $2.5$PN terms.
+We include both conservative 11PN and 22PN terms, and radiative 2.52.5PN terms.
 The latter term generates the GW emission and leads to the decrease in binary SMA ($\dot{a}_{\rm GW}$) and eccentricity only due to GWs.
@@
 Our approach is similar to the one outlined in Sect.
-6.2 of Liptai & Price (2019), except that we use a predictor-corrector approach instead of implementing the implicit kick-drift-kick one.
-In the initial setup the disk extended from $2a$ to $10a$.
-However during the first 1000 binary orbits, the cavity becomes eccentric and the inner edge increases to $\sim 3.5a$.
-We also run two other simulations with the same setup.
-First, without the dissipative $2.5$PN term to consider the binary evolution only due to the presence of the CBD.
-We run it for $1000$ orbits to see any appreciable changes in the orbital quantities.
-This allows us to infer the SMA evolution rate only due to the gaseous disk ($\dot{a}_{\rm gas}$) and therefore to disentangle its effect from the putative GW-gas cross term considered later.
-We then also run a simulation without the CBD in order to self-consistently obtain the evolution due to the PN terms only.
-This allows us to mitigate the numerical error in the integration when we compute the difference between the simulation with concurrent evolution under GWs and gas, and the one with PN terms only.
+6.2 of Liptai & Price (2019), except that we use a predictor-corrector approach instead of implementing the implicit kick-drift-kick one (see Franchini et al.
+(2024) for more details).
+In order to measure the effect of the gas contribution in the absence of GW emission, we run a simulation with the CBD, but without the 2.5PN dissipative term in the binary orbital motion.
+We refer to this simulation with the term “gas+PN2” and we perform it only at mid-resolution, i.e.
+with $\Delta x[3a]={1.07~r_{s}}$.
+We run this simulation for the same time as the gas+PN2.5 run to see any appreciable changes in the orbital quantities due to the sole interaction with the gaseous disk.
+This allows us to infer the gas torques and gas-induced orbital as well as binary precession dephasings without the effects introduced by GWs dissipation.
+Note that we extrapolate the mean torque results of the gas+PN2 simulation over the same integrated time down to $48.9~r_{\rm s}$ since evolving the binary under the mere effect of gas to such small separations is currently computationally prohibitive.
+Similarly, we compare the binary evolution between different setups over the same integrated time to compute the dephasings.
+We then also run two simulations without the CBD in order to self-consistently obtain the binary evolution driven only by the PN terms.
+We run one simulation including all the PN corrections, including the dissipative GW term, and another simulation with only the 1+2 PN corrections.
+We refer to the first and second simulation with the label “PN2.5” and “PN2” respectively.
+These simulations allow us to isolate the effects of GW emission and of the interaction with the disk, and to mitigate numerical errors in the integration when we compute the difference between the simulation with gas and their non-gaseous counterparts.
 2.1 Post-processing analysis
-The CBD typically affects the binary evolution by exerting both a gravitational torque ($T_{\rm grav}$) and an accretion torque ($T_{\rm acc}$).
+The CBD affects the binary evolution by exerting both a gravitational torque ($T_{\rm grav}$) and an accretion torque ($T_{\rm acc}$).
 The first, just due to gravity, is essentially driven by any asymmetry in the gaseous flow while the latter is instead induced by the accretion of gas particles onto either MBHs.
@@
 2021 for detailed calculations).
-We can then express the overall gas effect in terms of a single simulation-calibrated parameter (Garg et al., 2022; Duffell et al., 2024)
+We can then express the overall gas effect in terms of a single dimension-less simulation-calibrated parameter (similar to the accretion eigenvalue mentioned in Duffell et al. 2024)
 $\xi=\frac{T_{\rm grav}+T_{\rm acc}}{\dot{M}a^{2}\Omega},$ (1)
-where $\dot{M}\equiv 0.02~{}{\rm f}_{\rm Edd}(M/10^{6}{\rm M}_{\sun}){({\rm M}_{\sun}/\rm yr)}$ is the accretion rate onto the binary for our assumed $\epsilon=0.1$ radiative efficiency, $\Omega\equiv\sqrt{GM/a^{3}}$ is the binary orbital angular frequency, and $\dot{M}a^{2}\Omega$ is the viscous torque (Lin & Papaloizou, 1986).
-$\xi$ depends sensitively on the binary and disk parameters, in particular on the binary mass ratio $q$, disk shape and temperature, and may also depends upon the assumed equation of state.
-Previous sub-pc fixed binary orbit 2D Newtonian simulation predicted $|\xi|\sim 3$ (Dittmann & Ryan, 2022) for $H/R\sim 0.03$.
-We can therefore write down the SMA evolution rate due to gas torques for a quasi-circular binary as (Garg et al., 2024a)
-$\dot{a}_{\rm gas}=8\xi\frac{\dot{M}_{z}}{M_{z}}a=\xi{\rm f}_{\rm Edd}\frac{4}{25{\rm Myr}}a,$ (2)
-where we infer both $\xi$ and ${\rm f}_{\rm Edd}$ from gas-only simulation to properly isolate any GW-gas coupling in the total SMA rate in the GW+gas simulation.
-Moreover, since the evolution only due to gas torques changes the SMA by a small quantity even after $1000$ orbits, we extrapolate the results to smaller separations, i.e.
-until $30r_{s}$.
-When the MBHB is dynamically evolving under both GWs and gas, there can be a coupled evolution that manifests in the form of a cross term $\dot{a}_{\rm GWgas}$, whose exact functional form is unknown.
-Therefore, the total SMA evolution rate can be written down as
-$\dot{a}_{\rm GW+gas}=\dot{a}_{\rm GW}+\dot{a}_{\rm gas}+\dot{a}_{\rm GWgas},$ (3)
-where both $\dot{a}_{\rm gas}$ and $\dot{a}_{\rm GWgas}$ can be thought of as perturbative corrections to $\dot{a}_{\rm GW}$ in the final years of in-spiral.
-Assuming the stationary phase approximation (Cutler & Flanagan, 1994), where gas affects the phase but not the amplitude of the GW, we compute the dephasing in the time-domain GW waveform, caused by the presence of the CBD as
-$\delta\phi_{\rm gas}=\Phi(\dot{a}_{\rm GW+gas})-\Phi(\dot{a}_{\rm GW}),$ (4)
-where negative (positive) dephasing imply that gas speed-up (slow-down) the binary in-spiral.
-Here $\Phi(\dot{a})$ is the total accumulated GW phase for a given evolution rate $\dot{a}$ between an initial ($a_{i}$) and a final ($a_{f}$) SMAs, and it can be expressed in radians as
-$\Phi(\dot{a})=2\pi\int_{a_{\rm i}}^{a_{\rm f}}{\rm d}a\frac{f}{\dot{a}},$ (5)
-where $f\equiv\Omega/\pi$ is the GW frequency, that is twice the orbital frequency ($\Omega/2\pi$), under the assumption of quasi-circular binaries.
+Here $\Omega\equiv\sqrt{GM/a^{3}}$ is the binary orbital angular frequency, and $\dot{M}a^{2}\Omega$ is simply the normalization commonly used in the literature (see, e.g.
+Duffell et al., 2024).
+$\xi$ depends sensitively on the binary and disk parameters, in particular on the binary mass ratio $q$, disk shape and temperature, and may also depend upon the assumed equation of state.
+Previous 2D Newtonian simulations, featuring sub-pc fixed binary orbit, predicted $|\xi|\lesssim 2$ (Dittmann & Ryan, 2022) for $H/R\sim 0.03$, although for a higher kinematic viscosity value than the one we simulate in this work.
+Note that, due to our live binary setup, we can directly compute the change in SMA and eccentricity from the positions and velocities of the binary components, as well as the accretion rate.
+We can therefore measure $\xi$ and ${\rm f}_{\rm Edd}$ independently for both gas+PN2.5 and gas+PN2 simulations.
+The inclusion of gas in the MBHB system induces perturbations in the evolution of the binary orbital phase.
+The difference in the number of binary cycles before the merger when it evolves in a gaseous environment is the result of a combination of environmental effects.
+Indeed the gas causes a different evolution of the binary semi-major axis, eccentricity and precession rate.
+Since we lack analytical prescriptions for each of these terms in the PN approximation, we can only directly measure the gas-induced orbital ($\delta\phi_{\rm orb}$) and precession ($\delta\phi_{\rm ecc}$) dephasings by comparing the simulations with and without the gaseous disk.
+We therefore compare the gas+PN2.5 and PN2.5 orbital phases to infer the global effect that the gaseous disk has on the number of binary cycles of a GW-driven MBHB.
+We then compare the gas+PN2 and the PN2 runs to better isolate the effect of gas on the binary precession rate without any perturbations to the binary induced by GW emission.
+Indeed in the gas+PN2 and the PN2 setups the binary semi-major axis and eccentricity remain almost constant with time.
+Since most of the dephasing is accumulated at larger separations, where the coupling with the disk is stronger, we align the relevant pairs of simulations as much as possible, given the temporal resolution limitation in our simulations, in phase (set to $\pi$) at the final time.
+The ideal approach would be to align the phases exactly at merger but evolving our systems further would significantly increase the computational cost without adding significant information on the dephasing accumulated at larger separations.
+We then measure $\delta\phi_{\rm orb}$ by comparing the orbital phases of the appropriate pairs of simulations at the initial time.
+We measure $\delta\phi_{\rm ecc}$ by comparing the argument of periapsis $\omega=\tan^{-1}(e_{y}/e_{x})$ of the appropriate pairs of simulations, where $e_{x}$ and $e_{y}$ are the $x$ and $y$ components of the binary eccentricity vector $\vec{e}$.
+Lastly, under the quasi-circular approximation, we can infer the GW dephasing ($\delta\phi_{\rm GW}$) as twice the orbital dephasing, i.e., $\delta\phi_{\rm GW}\approx 2\delta\phi_{\rm orb}^{\rm(GW)}$.
+We should in principle compute the dephasing in each eccentric GW harmonic, since our binary is eccentric to begin with.
+However, the scope here is to report a simple estimate of GW dephasing for intuition and the majority of the power would be anyway contained in the fundamental harmonic (i.e at twice the orbital frequency) for our choice of parameters (Peters & Mathews, 1963).
 In the next section, we show the results together with the interpretation of our simulations.
 3 Results
-We show disk morphologies and associated column densities at three SMAs – $a=50r_{s}$, $a=40r_{s}$, and $a=30r_{s}$ – in Fig.
-1 for the GW+gas simulation.111disk morphology for gas-only simulation is similar to the left panel of Fig.
+We show the disk morphology in terms of column density at three SMAs – $a=54.5~r_{s}$, $a=52~r_{s}$, and $a=49.5~r_{s}$ – in Fig.
+1 for our MR gas+PN2.5 simulation.
+[^2] The disk morphology for gas+PN2 simulation is similar to the left panel of Fig.
 1.
 The gas morphology is similar, as expected, to the simulations presented in Franchini et al.
-(2024).
+(2024) with the addition of short-lived mini-disks owing to our higher resolution.
 We can clearly see the over-density at the cavity edge, i.e.
 the lump (Shi et al., 2012), precessing around the binary.
-We measure the accretion rate in both our simulations and find a mean Eddington ratio of ${\rm f}_{\rm Edd}\approx 0.92$ for GW+gas simulation and ${\rm f}_{\rm Edd}\approx 1.20$ for the gas-only case.
-The slightly higher accretion rate in the gas-only case is a natural consequence of the fact that the binary does not start decoupling from the disk because of the lack of GW radiation, therefore the gas can flow inside the cavity and keep feeding the MBHB more effectively.
-Figure 1: Column density ($\Sigma$) plots at three SMAs: $50r_{s}$ (left panel), $40r_{s}$ (middle panel), and $30r_{s}$ (right panel) for the binary evolution under both GW and gas.
-Here $\Sigma$ varies between $\sim 10^{3.5}$-$10^{6}$ g/cm2.
-While both the binary (green dots) and the cavity shrinks with time, cavity is becoming more eccentric also.
-We then compute both gravitational and accretion torques directly from the simulations.
+We measure the accretion rate in both our simulations and find a mean Eddington ratio of $\bar{{\rm f}_{\rm Edd}}\approx 1.30$ for gas+PN2.5 simulation and $\bar{{\rm f}_{\rm Edd}}\approx 1.50$ for the gas+PN2 case, both consistent with the analytical expectation for the choice of disk mass ($5{\rm M}_{\sun}$) we made and assuming a steady-state $\alpha$-disk with the properties we chose (Frank et al., 2002).
+The slightly higher accretion rate in the gas+PN2 case is a natural consequence of the fact that the binary does not decouple from the disk because of the lack of GW radiation, therefore the gas can flow inside the cavity and keep feeding the MBHB more effectively.
+Figure 1: Column density ($\Sigma$) plots at three SMAs: $54.5~r_{s}$ (left panel), $52~r_{s}$ (middle panel), and $49.5~r_{s}$ (right panel) for the binary evolution under both GW and gas.
+Here $\Sigma$ varies between $\sim 10^{3}$-$10^{7}$ g/cm2.
+Both the binary (green dots) and the cavity shrink with time.
+Moreover, gas inflow inside the cavity creates short-lived mini-disks.
+We compute both gravitational and accretion torques directly from the simulations.
 We find that the magnitude of the accretion torque is $T_{\rm acc}\sim 10^{-2}T_{\rm grav}$.
 We show the evolution of the parameter $\xi$ in Fig.
-2 for GW+gas simulations as a function of SMA together with the average values of $\xi$ for both GW+gas ($\bar{\xi}_{\rm GW+gas}\approx-24$) and gas-only ($\bar{\xi}_{\rm gas}\approx-36$) cases.
-Unsurprisingly, the mean gas effect is stronger when the binary is only evolving due to the gas torque.
-The parameter $\xi$ is indeed decreasing in magnitude as the binary is in-spiraling further, since the gas effects are becoming weaker.
-Since the value of $\xi$ does not change significantly with time in the gas-only simulation, the approximation we make in the parametrization of Eq.
-(1) is reasonable.
-It is worth noticing that for the computed gas-induced dephasing $\delta\phi_{\rm gas}$ (see Section 3.1), most of the contribution is expected to originate from the early in-spiral phase, where the mean value of $\xi$ is higher.
+2 for gas+PN2.5 simulations as a function of SMA together with the average values of $\xi$ for both gas+PN2.5 ($\bar{\xi}_{\rm{gas+PN2.5}}=-19.5$) and gas+PN2 ($\bar{\xi}_{\rm{gas+PN2}}=-24.4$) cases.
+Negative torques are consistent with similar aspect ratio HD studies (Tiede et al., 2020; Heath & Nixon, 2020; Dittmann & Ryan, 2022; Franchini et al., 2022).
+Unsurprisingly, the mean gas effect is stronger when the binary is only evolving due to the gas torques.
+Since the value of $\xi$ oscillates around its average value over time, one may approximately infer it as a constant parameter within the range of separations explored in this work.
+Figure 2: Gas torques onto the binary in terms of $\xi$ as a function of the SMA for gas+PN2.5 simulation (light blue lines).
+We show average $\xi$ values for gas+PN2.5 (solid blue line; $\bar{\xi}_{\rm{gas+PN2.5}}\approx-19.5$) and gas+PN2 (dashed red line; $\bar{\xi}_{\rm gas+PN2}\approx-24.5$) cases, respectively.
+Note that we have not done any smoothing in plotting $\xi$ for the gas+PN2.5 run but only interpolation between snapshots.
 We also note that the high-frequency or sub-orbital fluctuations in Fig.
-2 around the mean value may be measurable on their own, as suggested by analytical studies (Zwick et al., 2021, 2024).
-However, we focus on the secular trend in this work.
-Figure 2: Gas torques onto the binary in terms of $\xi$ as a function of the SMA for GW+gas simulation (light blue lines).
-We show average $\xi$ values for GW+gas (solid blue line; $\bar{\xi}_{\rm GW+gas}\approx-24$) and gas-only (dashed red line; $\bar{\xi}_{\rm gas}\approx-36$) cases, respectively.
-Once we subtract both $\dot{a}_{\rm GW}$ and $\dot{a}_{\rm gas}$ from the total SMA $\dot{a}_{\rm GW+gas}$ rate inferred from the full GW+gas simulation, we find a non-negligible residual, which is the GW-gas coupling term introduced in Eq.
-(3).
-We propose to express this cross-term simply as
-$\dot{a}_{\rm GWgas}\sim\zeta\dot{a}_{\rm GW}\dot{a}_{\rm gas},$ (6)
-where $\zeta$ is also a simulation-calibrated parameter and we show its evolution in Fig.
-3 as function of SMA.
-We find the average value of the cross-term to be $\approx-7.9\times 10^{-5}$.
-The value of $\zeta$ is also decreasing as the binary is shrinking, similar to the evolution of $\xi$ in Fig.
+2 around the mean value may be measurable on their own via GWs, as suggested by analytical studies (Zwick et al., 2022, 2024, 2025).
+However, recently, Copparoni et al.
+(2025) demonstrated, using realistic LISA data analysis, that these moderate fluctuations we find are not measurable in GWs, albeit for a much smaller mass ratio BHB compared to the one explored in our work.
+We show torque density maps averages over ten snapshots for the gas+PN2.5 (left panel) and gas+PN2 (middle panel) simulations together with their difference (right panel) in Fig.
+3.
+Similar to previous HD studies (see, e.g.
+Tiede et al., 2020; Franchini et al., 2023), we rotate gas particles about the binary’s center-of-mass such that the axis connecting the two BHs aligns in the two simulations to do a meaningful comparison between their gas distribution.
+We chose to average over ten orbits, as it is small enough that the binary is not significantly shrinking and large enough to remove instantaneous features.
+The first two panels show the contribution from both CBD and mini-disks to the gravitational torque.
+The gas in front of the BH tends to drag it (pink color) along while the gas behind it decelerates it (green color).
+The right panel clearly shows that the gravitational torque is larger in the gas+PN2 simulation as there are more regions where the torque is negative.
+This is again consistent with the estimate of the $\xi$ parameter in Fig.
 2.
-The almost constant evolution of $\zeta$, marginally weakened by the in-spiral, may imply that our assumption to express $\dot{a}_{\rm GWgas}\propto\dot{a}_{\rm GW}\dot{a}_{\rm gas}$ is not inaccurate.
-Figure 3: GW-gas coupling coefficient as a function of SMA (light blue lines) with its mean value $\bar{\zeta}=-7.9\times 10^{-5}$ (solid blue line).
-3.1 Gas-induced dephasing and LISA observability
-In this section, we compute the gas-induced dephasing $\delta\phi_{\rm gas}$ in Eq.
-4 in the source-frame between $=53.2~{}r_{s}$ and $=30~{}r_{s}$ with the initial eccentricity $=0.29$ by using $\dot{a}_{\rm GW}$ inferred from the simulations and then considering the total SMA evolution rate $\dot{a}_{\rm GW+gas}$ in three different ways.
-  • First of all directly from the GW+gas simulations, which leads to a dephasing $\delta\phi_{\rm gas}=-3.21$ radians.
-  • The second estimate of the dephasing is performed by neglecting the GW-gas cross term, i.e.
-assuming $\dot{a}_{\rm GW+gas}=\dot{a}_{\rm GW}+\dot{a}_{\rm gas}$, where $\dot{a}_{\rm gas}$ is computed from Eq.
-2 with the gas-only simulation-calibrated parameters $\xi=-36$ and ${\rm f}_{\rm Edd}=1.2$.
-This approach gives $\delta\phi_{\rm gas}=-0.21$ radians.
-  • Lastly, we include the cross-term using the parameterization in Eq.
-6 to have $\dot{a}_{\rm GW+gas}=\dot{a}_{\rm GW}+\dot{a}_{\rm gas}+\zeta\dot{a}_{\rm GW}\dot{a}_{\rm gas}$ with $\zeta=-7.9\times 10^{-5}$.
-This results in $\delta\phi_{\rm gas}=-3.63$ radians.
-The dephasing we obtain neglecting the GW-gas cross term (i.e.
-assuming $\zeta=0$) is an order of magnitude smaller than the value inferred with the non-zero cross term, so we caution that the GW-gas coupling might have a crucial role in the evolution of MBHBs in gaseous environments.
-Furthermore, the fact that the $\delta\phi_{\rm gas}$ computed directly from the GW+gas simulation is almost the same as the dephasing computed by parametrizing the gas effect through the (simulation-calibrated) $\xi$ and $\zeta$ parameters, provides a further consistency check of our model.
-The few percent difference between the two values might originate from the approximation made in assuming $\xi$ and $\zeta$ to be constants with time.
-We consider the evolution of our $10^{6}~{}{\rm M}_{\sun}$ total mass eccentric MBHB at redshift $z=1$ for one year of LISA observation.
-We therefore take an initial SMA of $a_{i}=40r^{z=1}_{s}$ and an initial eccentricity $e=0.2$, as the binary has partially circularized from its initial separation.
-We find the dephasing for ${\rm f}_{\rm Edd}=1.2$, $\xi=-36$, and $\zeta=-7.9\times 10^{-5}$ between $a_{i}=40r^{z=1}_{s}$ and $a_{f}=30r^{z=1}_{s}$ to be $\delta\phi_{\rm gas}=-1.11$ radians, equivalent to losing $\sim 0.2$ GW cycles over the expected $\sim 850$ cycles from a vacuum eccentric binary evolving just under GWs.
-Since LISA should observe this event with SNR1300similar-toabsent1300$\sim$ 1300∼ 1300 (Garg et al., 2024d) and we typically need a given dephasing to be higher than 8$\sim$ 8/∼ 8 /SNR (Derdzinski et al., 2021; Garg et al., 2022) to be detectable, we should be able to observe the computed $\delta\phi_{\rm gas}$.
-3.1.1 Analytical dephasing estimates
-Since we would ultimately like to obtain a parametrized model for the GW+gas inspiral phase, we now use the orbital-averaged GW-driven inspiral rate at $2.5$PN order from Peters (1964), i.e.
-$\dot{a}^{\rm Peters}_{\rm GW}$, for the GW only driven inspiral rate.
-Moreover, we also analytically construct the total SMA rate using the simulation-calibrated parameters $\xi$, ${\rm f}_{\rm Edd}$, and $\zeta$ to have
-$\dot{a}^{\rm analytical}_{\rm GW+gas}=\dot{a}^{\rm Peters}_{\rm GW}+\dot{a}_{\rm gas}+\zeta\dot{a}^{\rm Peters}_{\rm GW}\dot{a}_{\rm gas}.$ (7)
-If we repeat the dephasing computation done in § 3.1, then we get $\delta\phi_{\rm gas}=-3.75$ radians.
-This is a few percent higher than the value calculated using only the simulations data.
-One possible reason is the inclusion of 1PN and 2PN terms in the simulations which, despite being conservative terms, lead to the precession of the MBHB eccentric orbit effectively altering its in-spiral rate.
-Another source of uncertainty in the dephasing estimate might reside in numerical errors.
-We have however compared our GW-only simulation with a numerical integration of the full (i.e.
-up to 2.5 order) PN evolution of the binary using an 8th order Runge-Kutta integrator finding negligible differences within the separations range we explored.
-Even though we have only performed simulations for one set of binary-disk parameters, we can compute the expected dephasing for different values of the $\xi$ and $\zeta$ parameters, which reflect the use of different system parameters in the simulations, through the analytical prescription in Eq.
-7.
-In Fig.
-4, we show $\delta\phi_{\rm gas}$ for equal-mass circular MBHBs accreting at ${\rm f}_{\rm Edd}=1$ for a one-year LISA observation until respective ISCO with varying $\xi\in[-1,-100]$, $\zeta\in[0,2\times 10^{-4}]$, $M\in\{10^{5},10^{6}\}~{}{\rm M}_{\sun}$, and $z\in\{1,2\}$ to bracket reasonable parameters space.
-If we do not consider the GW-gas cross-term then the dephasing cannot be detectable by LISA for $M=10^{6}~{}{\rm M}_{\sun}$ MBHB at both redshifts for $|\xi|\lesssim 1$.
-The stronger dephasing with non-zero $\zeta$ again signifies the importance of considering the GW-gas cross-term in having observable gas effects on the GW waveform.
-Moreover, similar $\delta\phi_{\rm gas}$ for different combinations of mass and redshift is possibly due to the cross-term dependence on $\dot{a}^{\rm Peters}_{\rm GW}$.
-Figure 4: The gas-induced dephasing in the GW waveform, $\delta\phi_{\rm gas}$ in radians, over a one-year LISA observation until ISCO as a function of the simulation-calibrated fudge factors $\xi$ (represents gas torques) and $\zeta$ (represents GW-gas cross-term) for equal-mass MBHBs accreting at ${\rm f}_{\rm Edd}=1$.
-In all four panels, we vary $\xi$ from $-1$ to $-100$ and $\zeta$ from 00 to $2\times 10^{-4}$.
-In the top two panels, we consider systems at redshift $z=2$ and in the bottom two panels we assume $z=1$.
-In the left two panels, we study a MBHB with source-frame total mass $M=10^{5}~{}{\rm M}_{\sun}$ and in the right panels we have $M=10^{6}~{}{\rm M}_{\sun}$.
-We also draw cyan lines to indicate the regions of non-observable dephasing by LISA, i.e.
-$\delta\phi_{\rm gas}<8/$SNR.
-A perturbative correction to $\dot{a}_{\rm GW}$ with a functional form $\propto a^{\rm ng}\dot{a}_{\rm GW}$ manifests at $-n_{g}$ relative order in the GW phase (Garg et al., 2024a).
-We have $n_{g}=4$ and $n_{g}=1$ for $\dot{a}_{\rm gas}$ and $\dot{a}_{\rm GWgas}$, respectively.
-Hence, an instantaneous frequency-depending dephasing enters the overall binary phase at relative $-4$PN order due to evolution just under gas and at $-1$PN relative order due to a GW-gas coupling.
-Lastly, if we do not introduce a GW-gas coupling and try to explain the residual $(\dot{a}_{\rm GW+gas}-\dot{a}_{\rm GW})$ by using the model for $\dot{a}_{\rm gas}$ in Eq.
-2, then the mean value of the parameter $\xi$ is $\sim-360$.
-This implies that even a dephasing at $-4$PN relative order should be observable for our studied MBHB with or without concurrent eccentricity measurement (Garg et al., 2024a) in a realistic Bayesian data analysis setup, where all other binary parameters are also varying and a complete LISA response is considered.
-However, the GW-gas coupling is even more stronger than this, since it enters the GW phase at $-1$PN relative order, therefore causing a larger dephasing.
+Figure 3: 2D projected gravitational torque ($T_{\rm grav}$) distribution between 5a-5a to $5a$ in both axes averaged over $100$ snapshots between $100^{\rm th}$ and $110^{\rm th}$ orbits.
+We show results from the gas+PN2.5 run (left panel), the gas+PN2 run (middle panel), and their difference (right panel).
+The third panel clearly shows that the gas+PN2.5 run has slightly more positive torque than the gas+PN2 simulation.
+3.1 Gas-induced dephasings and LISA observability
+In this section, we consider different dephasings introduced at the end of § 2.1.
+We enumerate source-frame gas-induced dephasings that corresponds to elapsed physical time between $54.5~r_{s}$ and $48.9~r_{s}$ for the gas+PN2.5 simulation over 278 orbits or $\sim 1745.8$ rad orbital phase in Table 1.
+We have compared our PN2.5 simulation with a numerical integration of the full (i.e.
+up to 2.5 order) PN evolution of the binary using an 8th order Runge-Kutta integrator finding mean relative error of $\lesssim 0.1\%$ within the separations range we explored.
+This illustrates that our orbital parameters are well measured in different simulations.
+  $\rm\bf Dephasing$}   $\bf Value[rad]$}
+  $\delta\phi_{\rm orb}^{\rm(GW)}$   0.007-0.007
+  $\delta\phi_{\rm ecc}^{\rm(GW)}$   0.012-0.012
+  $\delta\phi_{\rm orb}^{\rm(NoGW)}$   0.014-0.014
+  $\delta\phi_{\rm ecc}^{\rm(NoGW)}$   0.010-0.010
+  $\delta\phi_{\rm GW}$   0.014-0.014
+Table 1: Dephasings measured from our simulations: gas-induced orbital $\delta\phi_{\rm orb}^{\rm(GW)}$ and precession $\delta\phi_{\rm ecc}^{\rm(GW)}$ dephasings between gas+PN2.5 vs PN2.5 simulations and same dephasings but between gas+PN2 and PN2 runs ($\delta\phi_{\rm orb}^{\rm(NoGW)},\delta\phi_{\rm ecc}^{\rm(NoGW)}$).
+We also infer gas-induced GW dephasing by doubling $\delta\phi_{\rm orb}^{\rm(GW)}$.
+The comparison between the orbital phase evolution in the gas+PN2.5 and PN2.5 simulations, $\delta\phi_{\rm orb}^{\rm(GW)}$, allows us to quantify the overall contribution of the different perturbations induced by the gas disk in the GW-driven binary evolution, i.e.
+its effect on $\dot{a}$, $\dot{e}$, and the precession rate.
+We can directly measure the dephasing due to the different precession rates $\delta\phi_{\rm ecc}^{\rm(GW)}$ directly from the simulations.
+However this is still not completely independent from other changes in the orbital parameters.
+Moreover, GWs radiation further complicates the picture as it changes the gas morphology, which can further contribute to $\delta\phi_{\rm orb}^{\rm(GW)}$ in a non-linear manner.
+Comparing the gas+PN2 and the PN2 runs can alleviate some of these issues as these runs both neglect the GW-driven fast inspiral, allowing us to measure the dephasing due to disk-induced binary precession $\delta\phi_{\rm ecc}^{\rm(NoGW)}$ in a system where the semi-major axis and eccentricity of the binary do not significantly evolve with time.
+The results presented in Table 1 show that the orbital dephasing measured in the gas+PN2.5 simulation (first row) is smaller than the same dephasing inferred in absence of GW emission (third row).
+This is consistent with the effect of the gas becoming weaker as the binary inspirals.
+The estimate of dephasing due to the binary precession in the second row is affected by the change in binary eccentricity induced by GW emission and is therefore not directly comparable to the term in the first row.
+The precession-induced dephasing calculated from the NoGW simulations (fourth row) is more similar to the term in the first row.
+This seems to indicate that the main contributor to the binary orbital dephasing is the disk-induced precession of the binary eccentricity vector.
+However we restrain to draw such a strong conclusion as longer simulations together with a larger exploration of the parameter space is needed to understand the role that non-linear effects play in the ultimate orbital dephasing in GWs.
+Furthermore, we note that all the dephasing inferred from the simulations are subjected to the particles noise and to the precise alignment of the orbital phase and argument of periapsis at the end of the simulations.
+If we compute the gas-induced GW dephasing analytically using only the gas torques by linearly adding SMA rates due to GWs and gas (Garg et al., 2022; Dittmann et al., 2023; Duffell et al., 2024) then we get $\delta\phi_{\rm GW}^{(\xi)}=-0.095$ rad, which is a factor of $7$ higher than our direct estimate of $\delta\phi_{\rm GW}=-0.014$ rad.
+This difference could be due to i) the inadequacy of the analytical prescription, ii) the approximation made by simply linearly adding two terms (i.e.
+GW-driven and gas-driven SMA change) that mutually affect each other and iii) different numerical approaches.
+Follow-up simulations might shed more light onto the nature of this discrepancy.
+We note that our estimate of precession-induced dephasing of $\delta\phi_{\rm ecc}^{\rm(GW)}=-0.012$ rad over 278 orbits is in magnitude comparable to the analytical prescription by Tiede et al.
+(2024) that yields dephasing of $\sim 0.010$ rad.
+This indicates that the choice of numerical method has a limited effect on the precession-induced dephasing.
+[…]
 4 Discussion and Conclusion
-We studied the interaction of an equal mass MBHB with its surrounding CBD during the late in-spiral stage, with and without concurrent GW emission, using 3D hydrodynamical simulations with a live binary.
+We studied the interaction of an equal mass MBHB with its surrounding geometrically thin ($H/R=0.03$) CBD during the late in-spiral stage, with and without concurrent GW emission, using 3D hydrodynamical simulations with a live binary.
 This approach provided us with the first direct measurement of how surrounding gas torques the binary when its in-spiral is already governed by GW emission, by means of the estimate of the $\xi$ parameter in Eq.
 (1).
-We find that $\xi$ is $\mathcal{O}(10)$ stronger than the previously explored scale-free/sub-pc regime.
-However, we caution that the comparison may not be fair as previous simulations carried out for the larger separation regime are predominantly 2D, assume a fixed binary orbit modelled under Newtonian dynamics, and compute the effect of the energy and angular momentum loss by GW radiation in post-processing (Tang et al., 2018).
+We find that $\xi$ is $\mathcal{O}(10)$ stronger than in some of the previously explored scale-free/sub-pc regime.
+However, we caution that the comparison may not be fair as previous simulations carried out for the larger separation regime are predominantly 2D, assume a fixed binary orbit modeled under Newtonian dynamics, explore larger values of the viscosity parameter $\nu$ (Dittmann & Ryan, 2022), and compute the effect of the energy and angular momentum loss by GW radiation in post-processing (Tang et al., 2018).
+Note that if we compared our value of $\xi$ in the gas+PN2 simulation with the results presented in Tiede et al.
+(2025) with the same value of viscosity that we have and the highest resolution they explored, we find the difference to be a factor of two.
+We also find that the measured time-averaged torque $\xi$ becomes weaker with higher resolutions (see Fig.
+A1), which is the opposite trend to the one found by Tiede et al.
+(2025).
+The different behavior might be due to our 3D live-orbit PN treatment with respect to their 2D fixed-orbit Newtonian simulations and to the different Mach number, i.e.
+$\mathcal{M}\approx 33$, employed in our simulations rather than their $\mathcal{M}=40$.
 In particular, Duffell et al.
@@
 (2023) argued that fixing the binary orbit leads again to a different gravitational torque.
-Moreover, we find that, consistently with our expectations, $\xi$ is slightly weaker for GW+gas simulations ($\xi_{\rm GW+gas}\sim-24$) with respect to gas-only study ($\xi_{\rm GW+gas}\sim-36$), as shown in Fig.
+Moreover, we find that, consistently with our expectations, $\xi$ is slightly weaker for gas+PN2.5 simulations ($\bar{\xi}_{\rm{gas+PN2.5}}\sim-19.5$) with respect to gas+PN2 study ($\bar{\xi}_{\rm{gas+PN2}}\sim-24.4$), as shown in Fig.
 2.
-This is expected since the binary is decoupling from the gas in the GW+gas simulation and therefore the effect of gas weakens with time.
-Most importantly, we show that our results can be understood by introducing a novel coupling term between the energy loss by GWs and that due to gas in the orbital in-spiral rate, which we model through the product between the rate of change of the SMA due to GWs only and that due to gas only.
-The amplitude of this coupling term is pinpointed by a newly introduced simulation-calibrated parameter $\zeta$, whose evolution is shown in Fig.
-3 with an average value $\sim-7.9\times 10^{-5}$.
-This implies richer gas features in the GW waveform than previously expected.
-While gas torques still induce a correction to the GW phase at the usual $-4$PN relative order (Garg et al., 2024a), our modeling of GW-gas cross-term enters the phase at $-1$PN relative order, leading to stronger phase shift by $\mathcal{O}(10)$ and higher likelihood of detection of gas perturbations with LISA.
-The $-1$PN relative order term may also allow us to break the degeneracy with concurrent eccentricity measurement using LISA (Garg et al., 2024a).
-Moreover, having a GW-gas coupling may change the expectation how the binary parameters in the LISA band are correlated with their values at sub-pc scales (Garg et al., 2024c).
-Lastly, ignoring gas effects while modeling MBHBs may lead to stronger apparent violations of general relativity than previously anticipated (Garg et al., 2024b) due to the new $-1$PN relative order term.
-In this work, we studied one set of binary-disk parameters while in the future we plan on exploring the parameter space to understand how both $\xi$ and $\zeta$ change.
-This exercise may also allow us to analytically model both simulation-calibrated parameters $\xi$ and $\zeta$.
-For instance, if the binary is embedded in a relatively thick disk ($H/R=0.1$), we can expect to also have a significant contribution from mini-disks (i.e.
-disks around each binary component) that will introduce their own likely positive torque measurement (D’Orazio & Duffell, 2021; Franchini et al., 2022; Siwek et al., 2023; Tiede & D’Orazio, 2023) and probably a different sign for $\zeta$.
-Moreover, different numerical studies may also allow to check the robustness of our phenomenological model.
-Lastly, we also find that, in the gas-only simulation, the eccentricity is linearly decreasing with the SMA.
-This needs to be further investigated in both sub-pc and near-merger regimes, and we plan to do so in the future.
+This is expected since the binary is decoupling from the gas in the gas+PN2.5 simulation and therefore the effect of gas weakens with time.
 It is to be noted that we employ PN corrections as an approximation since we do not run general relativistic magnetohydrodynamical (GRMHD) simulations.
-Current GRMHD simulations only study the binary evolution just a few days before merger (Gutiérrez et al., 2022; Avara et al., 2023) to integrate only a few orbits because of the prohibitive computational cost.
-Therefore, since the majority of the gas-driven effects on the binary in-spiral occurs at separations $a\gtrsim 30~{}r_{s}$, our approach is currently the best available method to investigate the orbital dephasing due to the presence of the gas and where PN corrections are may be adequate.
-A possible caveat in our work is that our binary is moderately eccentric ($\sim 0.3$) just a few years before the merger, which will require the eccentricity to be extremely high when GWs take over at milli-pc scales.
-However, since $e\sim 0.3$ arises naturally from our initial condition requiring a steady-state disk before setting the physical scale of $a\sim 53r_{s}$, the only truly realistic way to initialize the system is to evolve the binary starting from a much wider separation.
+Current GRMHD simulations only study the binary evolution just a few days before merger (Gutiérrez et al., 2022; Avara et al., 2024) to integrate only a few orbits because of the prohibitive computational cost.
+Therefore, since the majority of the gas-driven effects on the binary in-spiral occurs at separations $a\gtrsim 48.9~r_{s}$, our approach is currently the best available method to investigate the orbital dephasing due to the presence of the gas and where PN corrections may be adequate.
+A caveat in this work is that we assume that the gas morphology will look the same at our initial separation for all our gas+PN2.5 and gas+PN2 simulations.
+In principle, one would need to start the binary at sufficiently large separations, approaching parsecs, that the GW radiation is completely negligible, and quantify the difference in the morphology of the gas distribution while the binary shrinks as the GW radiation gradually ensues relative to a case in which it is neglected.
+This is currently not possible due to the prohibitive computational costs.
+If anything, by starting the binary with the same initial condition in the gas+PN2.5 and gas+PN2 simulation at a separation at which GW radiation is already taking place, we are erring on the side of underestimating the back-reaction of the gas to the GW emission, which translates into a conservative estimate of the cross-term.
+Another possible caveat in our work is that our binary is moderately eccentric ($\sim 0.3$) just a few years before the merger, which will require the eccentricity to be extremely high when GWs take over at milli-pc scales.
+However, since $e\sim 0.3$ arises naturally from our initial condition requiring a steady-state disk before setting the physical scale of $a\sim 55r_{s}$, the only truly realistic way to initialize the system is to evolve the binary starting from a much wider separation.
 This, however, would increase the computational cost dramatically.
@@
 Furthermore, our work, being the first of its kind with PN dynamics and a live binary in 3D, while still assuming a simple isothermal equation of state, provides a starting point for future hydrodynamical studies with additional physics, including, for example, more realistic thermodynamics.
-Acknowledgements
-MG acknowledge support from the Swiss National Science Foundation (SNSF) under the grant 200020_192092.
+Acknowledgments
+MG acknowledge support from the Swiss National Science Foundation (SNSF) under the grant 200020_192092 and “GW-learn” grant agreement CRSII5 213497.
+AF acknowledges support provided by the “GW-learn” grant agreement CRSII5 213497 and the Tomalla Foundation.
+AL acknowledges support by PRIN MUR “2022935STW”.
+We thank the anonymous referee for helpful comments that improved this work.
+We futher thank Rohit Chandramouli, Andrea Derdzinski, Alexander Dittmann, Callum Fairbairn, Zoltan Haiman, Laura Sberna, Connar Rowan, Christopher Tiede, and Lorenz Zwick for useful discussions.
 The authors also acknowledge use of the NumPy (Harris et al., 2020) and Matplotlib (Hunter, 2007).
+Appendix A Resolution study
+Figure A1: Average torque value (blue cross) expressed in terms of $\bar{\xi}$ over initial 100 orbits of gas+PN2.5 simulations for three different resolutions: LR with $\Delta x[3a]={1.31~r_{s}}$, MR with $\Delta x[3a]={1.07~r_{s}}$, and HR with $\Delta x[3a]={0.66~r_{s}}$.
+In Fig.
+A1, we show the torque values $\bar{\xi}$ time-averaged over 100 initial binary orbits, or equivalently in the SMA range between $54.5~r_{s}$ and $52.6~r_{s}$, for three different resolutions of the gaseous disk.
+We quantify the resolution by measuring an equivalent inter-particle spacing $\Delta x[3a]$ evaluated at $R=3a$.
+We name the three simulations as: low-resolution (LR) with $\Delta x[3a]={1.31~r_{s}}$, mid-resolution (MR) with $\Delta x[3a]={1.07~r_{s}}$, and high-resolution (HR) with $\Delta x[3a]={0.66~r_{s}}$.
+Fig.
+A1 shows the value of $\bar{\xi}$ for the three resolutions.
+Since the values measured from our MR and HR runs are very similar, within $0.4\%$, we can conclude that the MR run is sufficiently converged and we therefore further evolve it until $278$ orbits to measure the dephasing induced by the interaction of the binary with the disk.
+We run the gas+PN2 run at the MR resolution of $\Delta x[3a]={1.07~r_{s}}$ for the same elapsed physical time in order to measure the differences between the two simulations (see Section § 3).
+Since the gas+PN2 MR setup has no fast inspiral due to GWs, we can reasonably assume it is also converged.